OpenStudy (anonymous):

show that if u and v are differentiable functions of x and u> ​​0, then the y = u ^ ---> y '= u ^ v (v' * ln (v * u '/ u) ?

zepdrix (zepdrix):

was that supposed to be y = u ^ v `implies` other stuff?

OpenStudy (anonymous):

\[\large y=u^v~~\iff~~\ln y=v\ln u\] Differentiate both sides with respect to \(x\): \[\frac{d}{dx}\ln y=\frac{d}{dx}v\ln u\\ \frac{1}{y}y'=\frac{d}{dx}[v]\ln u+v\frac{d}{dx}[\ln u]\\ \frac{1}{y}y'=v'\ln u+v\frac{1}{u}u'\\ ~~~~~~~~~~~~\vdots\]